# Explicit Mutual Information Maximization for Self-Supervised Learning

### v2

@@ I Introduction
Self-supervised learning (SSL) is a powerful technique aims to learn task-agnostic representations without relying on annotated training data. SSL methods commonly utilize pretext tasks to learn representation-extracting models, which can be generally categorized into generative-based and contrastive-based ones [1, 2]. In generative-based methods, the pretext tasks are designed to predict or reconstruct the input data [3, 4, 5]. In comparison, contrastive-based methods construct pretext tasks in the embedding space, and learn a representation-extracting model by discriminating between positive and negative pairs in the embedding space.
@@
Mutual information (MI) is an information-theoretic measure that can capture non-linear statistical dependencies between random variables and thereby serve as a measure of true dependence [17]. Theoretically, under the framework of Siamese networks based SSL, MI is an optimal metric to measure the dependence between the embeddings of the Siamese networks. However, directly constructing an objective based on MI is challenging as it relies on the analytic expression of the data distribution. Consequently, many existing methods can be seen as approximate implementations of the MI maximization criterion. For example, the InfoNCE loss based methods [7] employ a noise contrastive estimation (NCE) of MI [18], which is shown to be a lower bound of MI [19]. Besides, it has been shown in [20] that the VICReg method [9] implements an approximation of the MI maximization criterion.
In this work, based on the invariance property of MI, we show that explicit MI maximization can be applied to SSL under a generic condition of the data distribution and derive a new objective for SSL. This new objective uses only second-order statistics but involves intractable determinant calculation of high-dimensional matrices. We propose reformulation and optimization strategies to make it adequate for stable and efficient end-to-end training. The main contributions are summarized as follows:
  • We provide a new perspective for SSL objective design from the invariance property of MI established in information theory, which shows that explicit MI optimization can be applied to SSL under a generic condition of the data distribution. Based on this result, an MI objective using only second-order statistics is derived.
@@ II Method
Figure 1: The proposed objective explicitly measures the MI $I(Z;Z^{\prime})$ between the embeddings $Z$ and $Z^{\prime}$ generated by two identical networks $f_{\omega}(\cdot)$ that are fed transformed versions of the same sample $S$. Maximizing the MI $I(Z;Z^{\prime})$ not only maximizes the dependency between the embeddings $Z$ and $Z^{\prime}$ by minimizing their joint entropy $H(Z,Z^{\prime})$, but also maximizes their marginal entropy $H(Z)$ and $H(Z^{\prime})$, respectively, which naturally avoids trivial constant solutions.
@@ II-A The Maximum Mutual Information Criterion
As illustrated in Figure 1, we consider directly maximizing the MI between $Z$ and $Z^{\prime}$, i.e. a maximum mutual information (MMI) criterion as
@@
The MI can be expressed as $I\left(Z;Z^{\prime}\right)=H(Z)+H(Z^{\prime})-H(Z,Z^{\prime})$, where $H(Z,Z^{\prime})$ is the joint entropy of $(Z,Z^{\prime})$, $H(Z)$ and $H(Z^{\prime})$ are the marginal entropy of $Z$ and $Z^{\prime}$, respectively. It is easy to see that the MMI criterion not only maximizes the dependency between the embeddings $Z$ and $Z^{\prime}$ by minimizing their joint entropy, but also maximizes the marginal entropy of $Z$ and $Z^{\prime}$, respectively, which promotes decorrelating the features in the embeddings $Z$ and $Z^{\prime}$ and naturally avoids the collapse problem of converging to trivial constant solutions.
However, MI is difficult to explicitly compute except for certain simple specific distributions. The next result shows that $I\left(Z;Z^{\prime}\right)$ can be explicitly computed based only on second-order statistics under a more generic distribution condition. In the following, we denote the covariance matrices of $Z,Z^{\prime}$ and $\tilde{Z}:=\left[Z^{T},Z^{\prime T}\right]^{T}$ by $C_{ZZ}$, $C_{Z^{\prime}Z^{\prime}}$ and $C_{\tilde{Z}\tilde{Z}}$, respectively.
@@
The proof is given in the Supplementary Material (SM) -B. This result is derived from the invariance property of MI [22] and the MI of multivariate Gaussian distribution. It relaxes the condition of data distribution required for explicit MI expression. It implies that we can compute the MI only based on second-order statistics even if the distributions of $Z$ and $Z^{\prime}$ are not Gaussian. To verify this result, we investigate the MI under a class of distributions, namely generalized Gaussian distribution (GGD). Generally, there does not appear to exist an agree on multivariate GGD, we use a definition as a particular case of the Kotz-type distribution [23, 24]
@@
The proof is given in SM -C. The above result implies that MI can be explicitly computed based only on second-order statistics for any distribution that can be marginally reparameterized as multivariate norm distribution under a homeomorphism condition.
@@ III Experiments
We follow the standard training protocol in the Solo-learn benchmark [25]. ResNet18 is used as the backbone for CIFAR-10/100 and ImageNet-100, whilst ResNet50 is used for ImageNet-1K. The results on CIFAR-10/100 and ImageNet-100 in Table I are compared with a batch size of 256. The experiments are conducted on a computing platform with three A100 GPUs, each with 40 GB of memory. On ImageNet-1K, it supports a maximum batch size of 1020 when using 16-bit precision (FP16). Therefore, on ImageNet-1K, our method is run with a precision of FP16 and a batch size of 1020. Meanwhile, to enhance stability and expedite the training process, we use gradient accumulation across four batches ($1020\times 4$) on ImageNet-1K. Detailed experiment settings are provided in the SM -E.
@@
Linear probing is a standard evaluation protocol in SSL, in which a linear classifier is trained on top of frozen representations of a pretrained model. Table I shows the linear probing results on CIFAR-10/100 and ImageNet-100. Following strictly the protocol of the Solo-learn benchmark [25], ResNet18 is used with the last fully connected layer replaced with a three-layer MLP interleaved with Batch Normalization (BN) and ReLU as projector. We keep the projector’s hidden dimension and output dimension as 2048. The modified backbone and projector act as encoder. For CIFAR, we adapt the architecture by removing the first max-pooling layer and modifying the first convolutional layer to fit the $32\times 32$ input size, in line with [25, 33, 7]. The results for other methods in Table I are sourced from the Solo-learn benchmark [25]. It can be seen that our method achieves competitive performance on all three datasets. Notably, on ImageNet-100, our method achieves the highest accuracy of 81.1%. Table II compares the top-1 accuracy of the methods on ImageNet-1K with different pretraining epochs. On ImageNet-1K, the results of our method are reported with a momentum encoder [15, 16]. As will be shown in Table III, our method can achieve better performance with a momentum encoder. It can be seen from Table II that our method can achieve competitive performance on ImageNet-1K, which is especially noteworthy under small training epochs.
@@
We evaluate the robustness of our method against batch size, following the same setting as Table I. We consider two variants of our method, one is standard and the other uses a momentum encoder [15, 16]. As shown in Table III, both our method and Barlow Twins reach a saturation point as batch size increases, which is consistent with the results in [8, 33]. Our method has better robustness across varying batch sizes, with a maximum accuracy variation of 1.6% on CIFAR-100 and 2.0% on ImageNet-100, respectively, while Barlow Twins shows 5.6% and 2.9% accuracy variation on the two datasets. For our method, using a momentum encoder helps to enhance performance across different batch sizes. More ablation studies can refer to the SM -D.

### v3

@@ I Introduction
Self-supervised learning (SSL) is a powerful technique aims to learn task-agnostic representations without relying on annotated training data. SSL methods commonly utilize pretext tasks to learn representation-extracting models, which can be generally categorized into generative-based and contrastive-based ones [1, 2]. In generative-based methods, the pretext tasks are designed to predict or reconstruct the input data [3, 4]. In comparison, contrastive-based methods construct pretext tasks in the embedding space, and learn a representation-extracting model by discriminating between positive and negative pairs in the embedding space.
@@
Mutual information (MI) is an information-theoretic measure that can capture non-linear statistical dependencies between random variables and thereby serve as a measure of true dependence [16]. Theoretically, under the framework of Siamese networks based SSL, MI is an optimal metric to measure the dependence between the embeddings of the Siamese networks. However, directly constructing an objective based on MI is challenging as it relies on the analytic expression of the data distribution. Consequently, many existing methods can be seen as approximate implementations of the MI maximization criterion. For example, the InfoNCE loss based methods [6] employ a noise contrastive estimation (NCE) of MI [17], which is shown to be a lower bound of MI [18]. Besides, it has been shown in [19] that the VICReg method [8] implements an approximation of the MI maximization criterion. In [20], the minimal coding length in lossy coding is used to construct a maximum entropy coding objective for SSL. Meanwhile, MI maximization is considered in [21] for SSL using a log-determinant approximation of the MI, which is simplified to derive a Euclidean distance-based objective.
In this work, based on the invariance property of MI, we show that the explicit MI maximization objective based on second-order statistics can be applied to SSL under a generic condition of the data distribution. This objective involves intractable determinant calculation of high-dimensional matrices, we propose reformulation and optimization strategies to make it adequate for stable and efficient end-to-end training. The main contributions are as follows:
  • We provide a new perspective for SSL objective design from the invariance property of MI established in information theory, which shows that explicit MI optimization based on second-order statistics can be applied to SSL under a generic condition of the data distribution.
@@ II Method
Figure 1: The MMI objective explicitly measures the MI $I(Z;Z^{\prime})$ between the embeddings $Z$ and $Z^{\prime}$ generated by two identical networks $f_{\omega}(\cdot)$ that are fed transformed versions of sample $S$. Maximizing $I(Z;Z^{\prime})$ not only maximizes the dependency between the embeddings $Z$ and $Z^{\prime}$ by minimizing their joint entropy $H(Z,Z^{\prime})$, but also maximizes their marginal entropy $H(Z)$ and $H(Z^{\prime})$, respectively, which naturally avoids trivial constant solutions.
@@ II-A The Maximum Mutual Information Criterion
As shown in Fig. 1, we consider a maximum mutual information (MMI) criterion as
@@
The MI can be expressed as $I\left(Z;Z^{\prime}\right)=H(Z)+H(Z^{\prime})-H(Z,Z^{\prime})$, where $H(Z,Z^{\prime})$ is the joint entropy of $(Z,Z^{\prime})$, $H(Z)$ and $H(Z^{\prime})$ are the marginal entropy of $Z$ and $Z^{\prime}$, respectively. The MMI criterion not only maximizes the dependency between the embeddings $Z$ and $Z^{\prime}$ by minimizing their joint entropy, but also maximizes the marginal entropy of $Z$ and $Z^{\prime}$, respectively, which promotes decorrelating the features in the embeddings $Z$ and $Z^{\prime}$ and naturally avoids the collapse problem of converging to trivial constant solutions.
However, MI is difficult to explicitly compute except for certain simple specific distributions. Accordingly, Gaussian distribution assumption is used to derive MI based objectives for SSL [21, 19]. The next result shows that $I\left(Z;Z^{\prime}\right)$ can be explicitly computed based only on second-order statistics under a more generic distribution condition. In the following, we denote the covariance matrices of $Z,Z^{\prime}$ and $\tilde{Z}:=\left[Z^{T},Z^{\prime T}\right]^{T}$ by $C_{ZZ}$, $C_{Z^{\prime}Z^{\prime}}$ and $C_{\tilde{Z}\tilde{Z}}$, respectively.
@@
The proof is given in the supplemental material (SM) [23]. This result is derived from the invariance property of MI [24] and the MI of multivariate Gaussian distribution. It relaxes the condition of data distribution required for explicit MI expression. It implies that we can compute the MI only based on second-order statistics even if the distributions of $Z$ and $Z^{\prime}$ are not Gaussian. To verify this result, we investigate the MI under a class of distributions, namely generalized Gaussian distribution (GGD). Generally, there does not appear to exist an agree on multivariate GGD, we use a definition as a particular case of the Kotz-type distribution [25, 26]
@@
The proof is given in the SM [23]. The above result implies that MI can be explicitly computed based only on second-order statistics for any distribution that can be marginally reparameterized as multivariate norm distribution under a homeomorphism condition.
@@ III Experiments
We follow the standard training protocol in Solo-learn benchmark [27]. ResNet18 is used as the backbone for CIFAR-10/100 and ImageNet-100, whilst ResNet50 is used for ImageNet-1K. The results on CIFAR-10/100 and ImageNet-100 in Table I are compared with a batch size of 256. The experiments are conducted on a computing platform with three A100 GPUs, each with 40 GB of memory. On ImageNet-1K, it supports a maximum batch size of 1020 when using 16-bit precision (FP16). Therefore, on ImageNet-1K, our method is run with a precision of FP16 and a batch size of 1020. Meanwhile, to enhance stability and expedite the training process, we use gradient accumulation across four batches ($1020\times 4$) on ImageNet-1K. Detailed experiment settings are provided in the SM [23].
@@
Linear probing is a standard evaluation protocol in SSL, in which a linear classifier is trained on top of frozen representations of a pretrained model. Table I shows the linear probing results on CIFAR-10/100 and ImageNet-100. Following strictly the protocol of the Solo-learn benchmark [27], ResNet18 is used with the last fully connected layer replaced with a three-layer MLP interleaved with Batch Normalization (BN) and ReLU as projector. We keep the projector’s hidden dimension and output dimension as 2048. The modified backbone and projector act as encoder. For CIFAR, we adapt the architecture by removing the first max-pooling layer and modifying the first convolutional layer to fit the $32\times 32$ input size, in line with [27, 6]. The results for other methods in Table I are sourced from the Solo-learn benchmark [27]. It can be seen that our method achieves competitive performance on all three datasets. Notably, on ImageNet-100, our method achieves the highest accuracy of 81.1%. Table II compares the top-1 accuracy of the methods on ImageNet-1K with different pretraining epochs. On ImageNet-1K, the results of our method are reported with a momentum encoder [14, 15]. As shown in Table III, our method achieves better performance with a momentum encoder. From Table II, our method achieves competitive performance on ImageNet-1K, which is especially noteworthy under small training epochs.
@@
We evaluate the robustness of our method against batch size, following the same setting as Table I. We consider two variants of our method, one is standard and the other uses a momentum encoder [14, 15]. As shown in Table III, both our method and Barlow Twins reach a saturation point as batch size increases, which is consistent with the results in [7]. Our method has better robustness across varying batch sizes, with a maximum accuracy variation of 1.6% on CIFAR-100 and 2.0% on ImageNet-100, respectively, while Barlow Twins shows 5.6% and 2.9% accuracy variation on the two datasets. For our method, using a momentum encoder helps to enhance performance across different batch sizes. More ablation studies can refer to the SM [23].